OpenStudy (anonymous):

weimen makes a square quilt that has a area of 100 ft it has 25 blocks each the same size what is the length of each side of the block a 2 b 4 c 6 d 10

OpenStudy (anonymous):

Just the answer of guiding you through it?

OpenStudy (anonymous):

Well.. you got offline... Fine let's see. We know we have a square of \( 100_{ft^2} \) divided to 25 equal sized blocks. That means... that: \[ \text{Block_Area} = \frac{\text{Quilt_Area}}{ \text{Number_of_blocks} } = \frac{100_{ft^2}}{25} = 4_{ft^2} \] Now.. assuming our blocks are squares, which means all sides are equal, the area of the block is: \[ \text{Block_Side}^2 = \text{Block_Area} \\ \text{Block_Side}^2 = 4_{ft^2} \\ \text{Block_Side} = \pm2_{ft} \] But since a side of a square can't be negative... \[ \text{Block_Side} \ge 0 \quad \implies \quad \text{Block_Side} \ne -2_{ft} \\ \text{Block_Side} = 2_{ft} \]